# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v1

@@ 4.4 Error Analysis and Explainability
The quantitative results highlight specific areas requiring deeper investigation, for instance, the poor performance of most detectors against EMOPortraits drobyshev2024emoportraits , particularly in P3 where T1 scores drop significantly. Similarly, CADMM dong2023implicit consistently underperforms on the TalkingHeadBench dataset, indicating a need to examine its limitations more closely.
We leverage explainability techniques like Grad-CAM Selvaraju_2019 to visualize the regions and features that detectors focus on for their predictions. We demonstrate an example for TALL xu2023tall since it achieves the overall best performance on our dataset. Other detectors can be found in the supplementary.
Failure Case Analysis: On the EMOPortraits dataset in P3-TALL xu2023tall ’s worst performing scenario (see Tab. 7), failure cases reveal that the model misclassifies both real and fake videos due to reliance on background cues. As shown in Fig. 5, real and fake videos are getting misclassified due to the model’s attention being disproportionately focused on the background, despite meaningful facial cues suggesting the correct label. This indicates a weakness in TALL’s generalization: it is overly influenced by background artifacts, likely due to training bias, and fails to localize attention to the face under combined identity and generator shifts.
@@
Success Case Analysis: In contrast to its failure cases, TALL demonstrates effective generalization in certain EMOPortraits drobyshev2024emoportraits videos under P3, as shown in Fig. 4. Here, the model accurately classifies a deepfake by focusing attention on the neck region—an area known to exhibit visual artifacts in EMOPortraits. Notably, this success occurs despite the fact that such artifacts are uncommon in other datasets used during training. This suggests that TALL xu2023tall has successfully learned to transfer knowledge from broader manipulation cues across datasets. The focused activation around the lower face and neck reflects meaningful model behavior and highlights its potential for cross-domain generalization in challenging settings.
@@
EMOPortraits drobyshev2024emoportraits 188 72 38.30%
Hallo2 cui2024hallo2 102 9 8.82%
LivePortrait guo2024liveportrait 168 5 2.98%
Table 8: Breakdown of high-confidence true positives (TPs) (prediction $\geq$ 0.9) in P1 that become false negatives (FNs) in P3 for TALL. A large drop from confident predictions in P1 to errors in P3 indicates poor generalization across generator shifts.
Tracking Generalization Failures from Confident Predictions: To better understand how detectors fail under distribution shifts, we track how high-confidence true positives (TPs) in P1 transition under the more challenging P3 setting. As shown in Tab. 8, for TALL—our best-performing detector—38.3% of EMOPortraits samples confidently classified in P1 (Prediction $\geq$ 0.9) are misclassified as false negatives in P3. This significant drop suggests TALL struggles to generalize when both identity and generator properties change. In contrast, TALL shows only a 2.98% drop from P1 to P3 on LivePortrait, highlighting that some generators pose far more severe generalization challenges. These results reinforce that P3 performance degradation is not random but rooted in generator-specific variation, with EMOPortraits emerging as a particularly valuable stress test for future detector evaluation.
Protocol-Specific Feature Reliance: To further understand detector behavior across generators, Fig. 6 presents Grad-CAM visualizations of TALL on samples from all six generators in Protocol 1, where we can study the bias of detectors to specific generator characteristics using the smallest train-test data distribution shift in P1. For generators such as AniPortrait (Audio/Video), Hallo, and LivePortrait, the model predominantly focuses on the central facial region, aligning well with human-interpretable deepfake cues. However, for Hallo2 and EMOPortraits, attention shifts away from facial features and toward areas surrounding the head. This may suggest one of two possibilities: either the facial features in these generators are realistic enough to fool the detector, or the model has learned to rely on peripheral artifacts (e.g., spatial warping in Hallo2 or neck distortions in EMOPortraits) as key detection signals. Such behavior indicates both progress in generator realism and potential over-reliance on non-facial cues in certain detector architectures.
@@ 4.5 Guidance for Future Research using TalkingHeadBench
Limitations of SOTA Detectors. Our analysis highlights that while models like TALL xu2023tall and DeepFake-Adapter shao2025deepfake demonstrate strong performance across most generators, there remains considerable room for improvement in achieving consistent excellence across all domains. Specifically, TALL achieves exceptional T1 scores on 3 out of 6 generators and maintains no worse than fair performance on the remaining ones, while DeepFake-Adapter reaches exceptional on 2 generators and stays within reasonable to fair on the rest. Notably, both models are only jointly exceptional on AniPortraitVideo wei2024aniportrait and Hallo xu2024hallo , suggesting current detectors may still struggle to generalize uniformly across diverse generation methods.
Open Challenges. While our analysis provides insights into generalization and robustness, several aspects of the TalkingHeadBench benchmark remain difficult for SOTA detectors. These include (i) handling challenging generators like EMOPortraits drobyshev2024emoportraits and Hallo2 cui2024hallo2 , where even the SOTA detectors reach only 70% and 88% T1 accuracy in P3; (ii) reducing performance degradation in P3 where both identity and generator properties differ between train-test splits, all detectors except DF-Adapter exhibit significant performance degradation from P1 to P3; (iii) overcoming model reliance overly on background artifacts—as demonstrated in Sec 4.4—which can cause the model to ignore relevant facial features and lead to misclassification. These persistent challenges highlight the need for future detectors to incorporate stronger domain adaptation, semantic attention, and modality-aware training to handle real-world deepfakes more reliably.
@@
(a) Performance: The best performing model across our analysis, TALL xu2023tall , reaches excellent performance in T1 ( $\>99\%$) in only 3/6 generators in P3, while reaching $\sim$90T1 for Hallo2 and LivePortrait and $\sim$80T1 for EMOPortraits. Future research should strive to reach the excellent performance ( $\>99\%$) in both T1 and more challenging T0.1 for all 6 detectors across P1, P2, and P3.
(b) Analysis: The TalkingHeadBench benchmark and protocols offer the granularity needed to pinpoint detector weaknesses and guide architectural improvements, training strategies, and robustness evaluation. Future research should prioritize reducing performance drop-offs in more challenging generators (e.g., EMOPortraits, Hallo2), exploring domain adaptation methods, and enhancing sensitivity to subtle facial artifacts, while avoiding overfitting to background or dataset-specific cues.
(c) Evolving Benchmark: A persistent challenge in deepfake detection is that benchmarks often lag behind the rapid progress of generative models, quickly becoming outdated. While the current TalkingHeadDetect repository on Hugging Face provides static train-test splits and evaluation protocols, in the near future, we aim to transform it into an evolving benchmark used by both the detector and generator communities. This will include (i) developing APIs that allow authors of new generators to evaluate the detectability of their models by submitting their codes, (ii) setting up dual leaderboards to track the progress of both generators and detectors, (iii) including generators and deepfake videos created by them whose detectability is lower than the ones present in the benchmark and (iv) removing generators that are easily detectable ($>99\%$ T0.1) a by majority of detectors, every six months. This evolving benchmark will encourage a healthy arms race that will accelerate innovation and improve the resilience of detection systems against newer generators.
@@ 5 Conclusions, Broad Impacts, and Limitations
We introduce TalkingHeadBench, a comprehensive benchmark utilizing contemporary talking-head generators to stress-test deepfake detectors under realistic distribution shifts. Our evaluations reveal critical gaps: even state-of-the-art detectors struggle with generalization across changes in identity and generator, particularly with challenging fakes like EMOPortraits drobyshev2024emoportraits , often due to over-reliance on non-facial artifacts. TalkingHeadBench will broadly impact the community by providing a standardized, challenging benchmark that reflects modern deepfake realism and complexity. We provide open access to the dataset, evaluation protocols, and baseline code. While we acknowledge one potential risk for the benchmark is that it could facilitate more sophisticated talking-head generation techniques, which could be potentially misused for deceptive or malicious purposes, our primary objective is to catalyze research into more robust, generalizable, and artifact-aware detection methods. The dataset only contains a finite number of videos and identities derived from FFHQ karras2019style and CelebV-HQ zhu2022celebv , which might not represent the diversity of real-world deepfakes that can be exhibited. In the future, we will set up leaderboards that allow performance comparison among newly developed generators and detectors. This benchmark serves as a crucial resource for developing trustworthy deepfake detection systems, essential for mitigating the increasing societal risks posed by manipulated media.
Appendices
All benchmarks and curated dataset for our TalkingHeadBench can be found in the following URL: https://huggingface.co/datasets/luchaoqi/TalkingHeadBench. TalkingHeadBench is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
@@
  • Details of Assets Used (Section A): Descriptions, sources, and licenses for source datasets, talking-head generators, and deepfake detectors utilized in our benchmark.
  • Detailed Experimental Setup (Section B): Dataset splits, data curation procedures, computational resources, and key training hyperparameters.
  • Detailed Per-Generator Artifact Observations (Section C): Specific types of visual artifacts commonly produced by each generator, noted during our manual curation process.
  • Experimental Results (Section D): Performance for commercial generator (MAGI-1).
  • Explainability Results (Section E): Grad-CAM visualizations for benchmarked detectors.
Appendix A Details of Assets Used
A.1 Source Datasets for Generation and Real Videos
@@
A.1.1 Source Images for Deepfake Generation
  • FFHQ (Flickr-Faces-HQ) [20]: This dataset served as the primary source for high-quality, diverse static portrait images used as the base for deepfake generation. It contains 70,000 high-resolution images of diverse human faces with diverse ages, ethnicities, and image backgrounds originally intended for GAN benchmarking. We utilized images from the 1024x1024 resolution subset. The dataset can be accessed through the following link: https://github.com/NVlabs/ffhq-dataset under Creative Commons BY-NC-SA 4.0 license by NVIDIA Corporation.
A.1.2 Source Driving Signals (Audio/Video) for Deepfake Generation
  • CelebV-HQ [55]: This large-scale video facial attributes dataset was used as the source for dynamic driving signals. It contains a variety of identities, ethnicities, expressions, and poses. Both video segments (for video-driven generators) and extracted audio tracks (for audio-driven generators, converted to .wav format) were utilized. The dataset can be accessed through the following link: https://github.com/CelebV-HQ/CelebV-HQ. The copyright information for this dataset is not explicitly mentioned.
A.1.3 Real Videos for Detector Training/Testing
  • FaceForensics++ (FF++) [35]: Original, unmanipulated video sequences were included in our set of real videos. We directly downloaded the dataset following the download script they provided, and finalized the number of our YouTube real videos from 1000 to 704 due to the unavailability of some of the videos on YouTube. We then renamed the downloaded audio to match the naming of the YouTube real videos. The renamed audio files are in our Hugging Face datasets (./audio/ff++). The original FF++ dataset can be accessed through the following link: https://github.com/ondyari/FaceForensics. The copyright information can be found at https://github.com/ondyari/FaceForensics/blob/master/LICENSE.
  • CelebV-HQ [55]: To diversify the identity in our real videos, we incorporated a distinct set of real videos from CelebV-HQ that were not used for driving signal extraction.
@@ A.1.3 Real Videos for Detector Training/Testing
A.2 Talking-Head Generators
@@ A.2.1 Hallo [49]
  • Key Artifacts Observed: See Section C.2.
A.2.2 Hallo2 [8]
@@
  • Key Artifacts Observed: See Section C.3.
A.2.3 AniPortrait [46]
@@
  • Key Artifacts Observed: See Section C.4.1 for audio-driven artifacts and C.4.2 for video-driven artifacts.
A.2.4 LivePortrait [14]
@@
  • Key Artifacts Observed: See Section C.1.
A.2.5 EMOPortraits [12]
@@
  • Key Artifacts Observed: See Section C.5.
A.2.6 MAGI-1 [38] (Commercial)
@@ A.3.1 CADDM [11]
A.3.2 TALL [50]
  • Description: TALL introduces a temporal-attentive localization and learning framework designed to exploit temporal inconsistencies in deepfake videos. By leveraging a dual-stream architecture that models both short-term and long-term temporal dependencies, TALL isolates manipulated frames and emphasizes temporal transitions typically ignored by frame-based detectors. A temporal attention mechanism further enhances frame-level representations by focusing on abrupt motion irregularities introduced by forgery generation processes. Moreover, the authors propose TALL-Swin, which integrates the TALL strategy with the Swin Transformer architecture. This combination leverages the Swin Transformer’s hierarchical feature representation and shifted window mechanism to effectively model both local and global dependencies within the thumbnail layout. Given there is no available model released on their original GitHub repository, we pretrained the model using FF++ based on the implementation version in DeepfakeBench [53].
@@ A.3.3 LipFD [25]
A.3.4 DeepFake-Adapter [40]
@@
Appendix B Detailed Experimental Setup
B.1 Dataset Generation, Splits, and Curation Details
B.1.1 Deepfake Video Generation
Approximately 500-600 videos were initially generated for each of the six open-source academic models. This was achieved by scripting random pairings of source images from FFHQ [20] with driving signals (audio or video) from distinct splits of CelebV-HQ [55].
B.1.2 Data Curation and Quality Control

### v2

@@ 4.5 Error Analysis and Explainability
The quantitative results highlight areas requiring deeper investigation, for instance, the poor performance of most detectors against EMOPortraits [drobyshev2024emoportraits], particularly in P3 where T1 scores drop significantly. Similarly, CADDM [dong2023implicit] underperforms consistently on TalkingHeadBench, indicating a need to examine its limitations more closely.
We leverage explainability techniques like Grad-CAM [Selvaraju_2019] to visualize the regions and features that detectors focus their predictions on. We demonstrate an example for TALL [xu2023tall] since it achieves the overall best performance on our dataset. Other detectors are in the supplementary.
Failure Case Analysis: On the EMOPortraits dataset in P3-TALL [xu2023tall]’s worst performing scenario (see Tab. 5(c)), failure cases reveal that the model misclassifies both real and fake videos due to reliance on background cues. As shown in Fig. 5, real and fake videos are getting misclassified due to the model’s attention being disproportionately focused on the background, despite meaningful facial cues suggesting the correct label. This indicates a weakness in TALL’s generalization: it is overly influenced by background artifacts, likely due to training bias, and fails to localize attention to the face under combined identity and generator shifts.
@@
Success Case Analysis: In contrast to its failures, TALL demonstrates effective generalization in certain EMOPortraits [drobyshev2024emoportraits] videos under P3, as shown in Fig. 4. The model accurately classifies a deepfake by focusing on the neck region, which is known to exhibit artifacts in EMOPortraits. Notably, this success occurs despite such artifacts being uncommon in other datasets used during training. This suggests TALL [xu2023tall] has transferred knowledge from broader manipulation cues across datasets. The focused activation around the lower face and neck reflects meaningful model behavior and highlights its potential for cross-domain generalization.
@@
EMOPortraits [drobyshev2024emoportraits] 188 72 38.30%
Hallo2 [cui2024hallo2] 102 9 8.82%
LivePortrait [guo2024liveportrait] 168 5 2.98%
Table 6: Breakdown of high-confidence true positives (TPs) ($\geq 0.9$) in P1 that become false negatives (FNs) in P3 for TALL. A large drop from confident predictions in P1 to errors in P3 indicates poor generalization across generator shifts.
Tracking Generalization Failures from Confident Predictions: To understand detector failures under distribution shifts, we track how high-confidence true positives (TPs) in P1 transition under P3. As shown in Tab. 6, for TALL—our best-performing detector—38.3% of EMOPortraits samples confidently classified in P1 (Prediction $\geq$ 0.9) are misclassified as false negatives in P3. This drop suggests TALL struggles to generalize when both identity and generator properties change. In contrast, TALL shows only a 2.98% drop from P1 to P3 on LivePortrait, highlighting that some generators pose far more severe generalization challenges. These results reinforce that P3 degradation is not random but rooted in generator-specific variation, with EMOPortraits emerging as a valuable stress test for detector evaluation.
Protocol-Specific Feature Reliance: To further understand detector behavior across generators, Fig. 6 presents Grad-CAM visualizations of TALL on samples from all six generators in Protocol 1, where we can study the bias of detectors to specific generator characteristics using the smallest train-test data distribution shift in P1. For generators such as AniPortrait (Audio/Video), Hallo, and LivePortrait, the model predominantly focuses on the central facial region, aligning well with human-interpretable deepfake cues. However, for Hallo2 and EMOPortraits, attention shifts away from facial features and toward areas surrounding the head. This may suggest: either the facial features in these generators are realistic enough to fool the detector, or the model has learned to rely on peripheral artifacts (e.g., spatial warping in Hallo2 or neck distortions in EMOPortraits) as key detection signals. Such behavior indicates both progress in generator realism and potential over-reliance on non-facial cues in certain detector architectures.
@@ 4.6 Guidance for Future Research using TalkingHeadBench
Limitations of SOTA Detectors. Our analysis shows that while models like TALL [xu2023tall] and DeepFake-Adapter [shao2025deepfake] perform strongly across most generators, there remains room for improvement in achieving consistent excellence across domains. Specifically, TALL achieves exceptional T1 scores on 3 of 6 generators and no worse than fair on the others, while DeepFake-Adapter reaches exceptional on 2 and stays within reasonable to fair elsewhere. Notably, both models are jointly exceptional only on AniPortraitVideo [wei2024aniportrait] and Hallo [xu2024hallo], suggesting detectors still struggle to generalize uniformly across diverse generation methods.
Open Challenges. While our analysis offers insights into robustness, several aspects of TalkingHeadBench remain difficult. These include (i) handling generators like EMOPortraits [drobyshev2024emoportraits] and Hallo2 [cui2024hallo2], where SOTA detectors reach only 70% and 88% T1 in P3; (ii) reducing degradation in P3 where both identity and generator differ between splits, as all detectors except DF-Adapter degrade significantly; (iii) overcoming reliance on background artifacts—as shown in Sec. 4.5—which causes models to ignore relevant facial features. These challenges highlight the need for detectors with stronger domain adaptation, semantic attention, and modality-aware training to handle real-world deepfakes more reliably.
@@
(a) Performance: The best model, TALL [xu2023tall], reaches excellent T1 ($99\%$) on only 3/6 generators in P3, while reaching $\sim$90% for Hallo2 and LivePortrait and $\sim$80% for EMOPortraits. Future research should strive for $99\%$ in both T1 and T0.1 across all 6 detectors and protocols.
(b) Analysis: TalkingHeadBench and its protocols offer the granularity to pinpoint weaknesses and guide improvements. Research should prioritize reducing drops in challenging generators (e.g., EMOPortraits, Hallo2), exploring domain adaptation, and enhancing sensitivity to subtle artifacts while avoiding overfitting to background cues.
(c) Evolving Benchmark: A persistent challenge is that benchmarks lag behind generative models, quickly becoming outdated. While TalkingHeadBench on Hugging Face provides static splits, we aim to evolve it for both detector and generator communities. This includes (i) APIs for authors to test detectability, (ii) dual leaderboards to track progress, (iii) adding generators with lower detectability, and (iv) removing generators easily detectable ($99\%$ T0.1) by most detectors every six months. This evolving benchmark will accelerate innovation and strengthen detection systems against newer generators.
@@ 5 Conclusions, Broad Impacts, and Limitations
We introduce TalkingHeadBench, a comprehensive benchmark utilizing contemporary talking-head generators to stress-test deepfake detectors under realistic distribution shifts. Our evaluations reveal critical gaps: even state-of-the-art detectors struggle with generalization across changes in identity and generator, particularly with challenging fakes like EMOPortraits [drobyshev2024emoportraits], often due to over-reliance on non-facial artifacts. TalkingHeadBench will broadly impact the community by providing a standardized, challenging benchmark that reflects modern deepfake realism and complexity. We will provide open access to the dataset, evaluation protocols, and baseline code.
While we acknowledge one potential risk for the benchmark is that it could facilitate more sophisticated talking-head generation techniques, which could be potentially misused for deceptive or malicious purposes, our primary objective is to catalyze research into more robust, generalizable, and artifact-aware detection methods.
The dataset only contains a finite number of videos and identities derived from FFHQ [karras2019style] and CelebV-HQ [zhu2022celebv], which might not represent the diversity of real-world deepfakes that can be exhibited. While we focus on creating high-quality forgeries, we acknowledge that our dataset can inherit the demographic imbalances from the public source datasets. We provide a detailed demographic audit in the Supplementary for transparency, and identify notable disparities across racial and age groups, while gender bias is minimal. In the future, we will set up leaderboards that allow performance comparison among newly developed generators and detectors. Additionally, we will conduct regular fairness audits, applying both group-based and individual fairness metrics to ensure robustness in our future leaderboard as new generators emerge. This benchmark serves as a crucial resource for developing trustworthy deepfake detection systems, essential for mitigating the increasing societal risks posed by manipulated media.
6 Acknowledgement
This research was supported in part by Lenovo Research (Morrisville, NC). We gratefully acknowledge the invaluable support and assistance of the members of the Mobile Technology Innovations Lab.
Appendix A Supplementary Materials
All benchmarks and curated dataset for our TalkingHeadBench are publicly released at https://anaxqx.github.io/talkingheadbench.github.io. TalkingHeadBench is released under a Creative Commons Attribution 4.0 License (CC BY 4.0).
@@ Appendix A Supplementary Materials
  • Details of Assets Used (Section B): Descriptions, sources, and licenses for source datasets, talking-head generators, and deepfake detectors utilized in our benchmark.
  • Detailed Experimental Setup (Section C): Dataset splits, data curation procedures, computational resources, and key training hyperparameters.
  • Detailed Per-Generator Artifact Observations (Section D): Specific types of visual artifacts commonly produced by each generator, noted during our manual curation process.
  • Experimental Results (Section E): Performance for commercial generator (MAGI-1) and academic generator (Hallo3).
  • Explainability Results (Section F): Grad-CAM visualizations for benchmarked detectors.
Appendix B Details of Assets Used
B.1 Source Datasets for Generation and Real Videos
@@
Source Images for Deepfake Generation
  • FFHQ (Flickr-Faces-HQ) [karras_style-based_2019]: This dataset served as the primary source for high-quality, diverse static portrait images used as the base for deepfake generation. It contains 70,000 high-resolution images of diverse human faces with diverse ages, ethnicities, and image backgrounds originally intended for GAN benchmarking. We utilized images from the 1024x1024 resolution subset. The dataset can be accessed through the following link: https://github.com/NVlabs/ffhq-dataset under Creative Commons BY-NC-SA 4.0 license by NVIDIA Corporation.
Source Driving Signals (Audio/Video) for Deepfake Generation
  • CelebV-HQ [zhu2022celebv]: This large-scale video facial attributes dataset was used as the source for dynamic driving signals. It contains a variety of identities, ethnicities, expressions, and poses. Both video segments (for video-driven generators) and extracted audio tracks (for audio-driven generators, converted to .wav format) were utilized. The dataset can be accessed through the following link: https://github.com/CelebV-HQ/CelebV-HQ. The copyright information for this dataset is not explicitly mentioned.
Real Videos for Detector Training/Testing
  • FaceForensics++ (FF++) [rossler2019faceforensics++]: Original, unmanipulated video sequences were included in our set of real videos. We directly downloaded the dataset following the download script they provided, and finalized the number of our YouTube real videos from 1000 to 704 due to the unavailability of some of the videos on YouTube. We then renamed the downloaded audio to match the naming of the YouTube real videos. The renamed audio files are in our Hugging Face datasets (./audio/ff++). The original FF++ dataset can be accessed through the following link: https://github.com/ondyari/FaceForensics. The copyright information can be found at https://github.com/ondyari/FaceForensics/blob/master/LICENSE.
  • CelebV-HQ [zhu2022celebv]: To diversify the identity in our real videos, we incorporated a distinct set of real videos from CelebV-HQ that were not used for driving signal extraction.
@@
B.2 Talking-Head Generators
@@
  • Key Artifacts Observed: See Section D.2.
Hallo2 [cui2024hallo2]
@@
  • Key Artifacts Observed: See Section D.3.
Hallo3 [cui2024hallo3]
  • Description: Hallo3 is an audio-driven portrait image animation model that uses a diffusion transformer backbone, with an identity reference network combining a causal 3D VAE plus stacked transformer layers, to generate highly dynamic, realistic video from static portraits. It improves over U-Net based generators by better handling non-frontal perspectives, immersive backgrounds, and motion dynamics, while preserving identity consistency even across viewpoint changes.
  • Modality: Audio-driven.
  • License: MIT License.
  • Key Artifacts Observed: Generally high quality; specific subtle artifacts, if any, are less pronounced or systematic compared to some open-source academic models.
AniPortrait [wei2024aniportrait]
@@
  • Key Artifacts Observed: See Section D.4 for audio-driven artifacts and D.4 for video-driven artifacts.
LivePortrait [guo2024liveportrait]
@@
  • Key Artifacts Observed: See Section D.1.
EMOPortraits [drobyshev2024emoportraits]
@@
  • Key Artifacts Observed: See Section D.5.
MAGI-1 [magi1] (Commercial)
@@ B.3 DeepFake Detectors
TALL [xu2023tall]
  • Description: TALL introduces a temporal-attentive localization and learning framework designed to exploit temporal inconsistencies in deepfake videos. By leveraging a dual-stream architecture that models both short-term and long-term temporal dependencies, TALL isolates manipulated frames and emphasizes temporal transitions typically ignored by frame-based detectors. A temporal attention mechanism further enhances frame-level representations by focusing on abrupt motion irregularities introduced by forgery generation processes. Moreover, the authors propose TALL-Swin, which integrates the TALL strategy with the Swin Transformer architecture. This combination leverages the Swin Transformer’s hierarchical feature representation and shifted window mechanism to effectively model both local and global dependencies within the thumbnail layout. Given there is no available model released on their original GitHub repository, we pretrained the model using FF++ based on the implementation version in DeepfakeBench [yan_deepfakebench_2023].
@@
DeepFake-Adapter [shao2025deepfake]
@@
AltFreezing [wang2023altfreezing]
  • Description: A video-based deepfake detector that leverages a spatiotemporal model. It employs a training strategy for 3D ConvNet video detectors that alternately freezes spatial- and temporal-related parameter groups to force learning of both artifact types, yielding stronger out-of-distribution generalization on face forgery videos.
  • Modality: Video-based (3D ConvNet / spatiotemporal).
  • License: MIT license.
MM-Det [song2025learningmultimodalforgeryrepresentation]
  • Description: A diffusion deepfake video detector that builds a Multi-Modal Forgery Representation (MMFR) using a large multi-modal model (LLaVA) and fuses it with a spatiotemporal backbone enhanced by In-and-Across Frame Attention (IAFA). Introduces the DVF dataset and reports SOTA on diffusion-generated videos.
  • Modality: Video-based (Diffusion).
  • License: Apache-2.0 license.
HiFi-Net [hifi_net_xiaoguo]
  • Description: An image forgery detection & localization framework with a hierarchical fine-grained formulation: multi-branch feature extractor plus dedicated localization and classification heads to capture subtle artifacts across manipulation types.
  • Modality: Image-based (detection + localization).
  • License: MIT license.
Appendix C Detailed Experimental Setup
C.1 Dataset Generation, Splits, and Curation Details
Deepfake Video Generation
Approximately 500-600 videos were initially generated for each of the six open-source academic models. This was achieved by scripting random pairings of source images from FFHQ [karras_style-based_2019] with driving signals (audio or video) from distinct splits of CelebV-HQ [zhu2022celebv].
Data Curation and Quality Control